Self-Study Problem 5.04 You have an opportunity to invest $4,800 today and receive $5,400 in three years. What would be the return on your investment if you accepted this opportunity? (Round answer to 2 decimal places, e.g. 52.75.) Return on investement
Answer #1

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

5400=4800*(1+r/100)^3

(5400/4800)^(1/3)=(1+r/100)

(1+r/100)=1.0400

r=1.0400-1

=4%(Approx).